# Scalable In-Context Q-Learning

### v2

@@ 4.6 Visualization Insights
We gain deep insights into the prompt construction process through t-SNE visualization on Cheetah-Vel and Ant-Dir tasks, as shown in Figure 6. We use a continuous color spectrum to indicate task similarity. For each task, we randomly sample 200 transitions and encode them into precise task representations using the pretrained world model. The initially entangled transitions are transformed into well-separated clusters in the prompt space, where points from different tasks are clearly distinguished and similar tasks are grouped more closely. Further, prompt representations of Cheetah-Vel form a clear rectilinear distribution from blue (low velocity) to red (high velocity), exactly aligning with the rectilinear spectrum of target velocities in a physical sense. Similarly, prompt representations of Ant-Dir follow a cyclic spectrum that matches the periodicity of angular directions in a physical sense. This interesting finding highlights S-ICQL’s ability to harness world modeling for distilling meaningful task-specific information from raw transitions, enabling precise task inference to facilitate in-context learning capacity.
@@ 5 Conclusions, Limitations, and Future Work
Though, the length of our prompt is the same as the sampled transitions, which may not be short enough, especially for long-horizon interactive problems. Future work can explore extracting task information with more compact structures, e.g., encoding an episode or a skill into a single token. Another promising step is to leverage natural language as a higher-level task prompt for ICLR, harnessing a wealth of knowledge within pretrained large language models (LLMs).
@@ Appendix E The Details of Baselines
  • UNICORN (Li et al., 2024b), integrates representative methods such as FOCAL (Li et al., 2021), CORRO (Yuan & Lu, 2022), and CSRO (Gao et al., 2023), this work proposes a unified information-theoretic framework that interprets these approaches as optimizing different approximation bounds of the mutual information between task variables and their latent representations. Building on the information bottleneck principle, it further derives a general and unified objective for task representation learning, facilitating the extraction of robust and transferable task representations. This method provides a unified information-theoretic perspective that summarizes and connects several mainstream context-based offline meta-RL approaches. As it integrates key ideas from representative COMRL algorithms, it can serve as a strong and generalizable baseline.
Mainstream ICRL algorithms predominantly rely on imitation learning paradigms, which tend to learn suboptimal behaviors when pretraining data is limited in size or quality. To ensure fair comparison, we standardize the dataset size and quality across methods during pretraining. In our task setting, AD struggles to extract effective policy improvement operators due to insufficient offline data. Therefore, we adopt an AD variant that incorporates a reward-based trajectory sorting mechanism from AT (Liu & Abbeel, 2023) to better distill policy improvements. Additionally, since UNICORN under the COMRL setting requires warm-up data for task representation inference, we align it with the ICRL setup by initializing task representations using randomly sampled trajectories and updating them after each episode.
@@ Appendix F Implementation Details of S-ICQL
Algorithm Hyperparameters. We evaluate the proposed S-ICQL algorithm on six environments: Darkroom, Push, Reach, Cheetah‑Vel, Walker‑Param, and Ant‑Dir. For all experiments, we use the Adam optimizer with a weight decay of 1e-4, gradient‑norm clipping at 10, an experience step of 4, and a batch size of 128. Table 8 summarizes the detailed hyperparameters of S-ICQL in each domain.
@@
Hyperparameters Darkroom Push Reach Cheetah-Vel Walker-Param Ant-Dir
Training steps 200000 400000 400000 400000 400000 400000
Learning rate 3e-4 1e-4 1e-4 3e-4 3e-4 3e-4
Prompt horizon $h$ 100 100 100 200 200 200
Embedding dim 32 128 128 128 128 128
Output layers num 1 2 2 2 2 2
Temperature parameter $\lambda$ 0.01 0.001 0.001 0.001 0.1 0.01
Expectile parameter $\omega$ 0.7 0.5 0.5 0.7 0.7 0.7
Soft updat $\alpha$ 0.005 0.005 0.005 0.005 0.005 e 0.005
Discount factor 0.99 0.99 0.99 0.99 0.99 0.99
Compute. We train our models on NVIDIA RTX3090 GPUs paired with an AMD EPYC 9654 CPU and 512GB of RAM. Pretraining the world model takes approximately 0.2–2 hours, while pretraining the causal transformer requires about 1–12 hours, depending on the environment’s complexity.
@@ Appendix G Hyperparameter Analysis
Appendix H Graphic Results for Ablation Study
Due to limited space, we only present tabular results in Sec. 4.2 in the main text. Here, we present the corresponding graphic results. Figure 9 shows the few-shot evaluation return curves of S-ICQL and its ablations on Mixed datasets.

### v3

@@ 4.6 Visualization Insights
We gain deep insights into the prompt construction process through t-SNE visualization on Cheetah-Vel and Ant-Dir tasks, as shown in Figure 8. We use a continuous color spectrum to indicate task similarity. For each task, we randomly sample 200 transitions and encode them into task representations using the pretrained world model. The initially entangled transitions are transformed into well-separated clusters in the prompt space, where points from different tasks are clearly distinguished and similar tasks are grouped more closely. Further, representations of Cheetah-Vel form a clear rectilinear distribution from blue (low velocity) to red (high velocity), exactly aligning with the rectilinear spectrum of target velocities in a physical sense. Similarly, representations of Ant-Dir follow a cyclic spectrum that matches the periodicity of angular directions in a physical sense. This finding highlights S-ICQL’s ability to harness world modeling to distill meaningful task-specific information from raw transitions, enabling precise task inference to facilitate in-context learning.
4.7 Additional Results on More Complex Environment
Figure 7: Few-shot evaluation return curves of all methods on Mixed datasets for complex tasks.
To further assess the scalability and robustness of S-ICQL in demanding scenarios, we extend our empirical evaluation to two challenging environments: (1) a sparse-reward, hard-exploration task (PickPlace) and (2) a high-dimensional control task (Humanoid-Dir), further increasing the diversity and difficulty of our evaluation. As shown in Figure 7 and Table 6, S-ICQL consistently outperforms all baselines, demonstrating robustness and generalization in these environments. Notably, the performance gap becomes even more pronounced in the high-dimensional setting, highlighting the particular strength of S-ICQL when dealing with large-scale problems, further underscoring its broader applicability.
Table 6: Few-shot converged returns of S-ICQL and baselines on Mixed datasets for complex tasks.
Method UNICORN DPT AD IDT DICP DIT IC-IQL S-ICQL
[…]
@@ 5 Conclusions, Limitations, and Future Work
Though, our prompt length matches the sampled transitions, which may be too long for long-horizon interactive problems. Future work can explore using more compact task representations, e.g., encoding an episode or a skill into a single token. Another promising step is to leverage natural language as a higher-level task prompt for ICLR. We also plan to utilize the dynamics decoder at test time to detect distributional shifts, facilitating adaptive learning through continual updates.
Acknowledgements
This work was supported in part by the National Natural Science Foundation of China (Nos. 62376122 and 72471109), in part by the National Key Research and Development Program of China (No. 2025YFA1016904), and in part by the National Natural Science Foundation of China (Nos. 92370132 and 72471109).
@@ Appendix C The Details of Environments
  • Humanoid-Dir: A high-dimensional humanoid agent must move in a specified direction while maintaining balance and coordinated whole-body control. Each task defines a target heading angle, and the reward is computed as the cosine similarity between the agent’s velocity vector and the unit vector of the target direction. Target directions are sampled uniformly from $0,2\pi]$.
@@
  • PickPlace: A Sawyer arm must pick up an object and place it at a target position sampled for each task. The object’s start pose and goal location vary across tasks. Rewards are sparse and reflect successful grasping and placement, making this a typical multi-step manipulation benchmark.
@@ Appendix E The Details of Baselines
  • UNICORN (Li et al., 2024b), integrates representative methods such as FOCAL (Li et al., 2021), CORRO (Yuan and Lu, 2022), and CSRO (Gao et al., 2023), this work proposes a unified information-theoretic framework that interprets these approaches as optimizing different approximation bounds of the mutual information between task variables and their latent representations. Building on the information bottleneck principle, it further derives a general and unified objective for task representation learning, facilitating the extraction of robust and transferable task representations. This method provides a unified information-theoretic perspective that summarizes and connects several mainstream context-based offline meta-RL approaches. As it integrates key ideas from representative COMRL algorithms, it can serve as a strong and generalizable baseline.
Mainstream ICRL algorithms predominantly rely on imitation learning paradigms, which tend to learn suboptimal behaviors when pretraining data is limited in size or quality. To ensure fair comparison, we standardize the dataset size and quality across methods during pretraining. In our task setting, AD struggles to extract effective policy improvement operators due to insufficient offline data. Therefore, we adopt an AD variant that incorporates a reward-based trajectory sorting mechanism from AT (Liu and Abbeel, 2023) to better distill policy improvements. Additionally, since UNICORN under the COMRL setting requires warm-up data for task representation inference, we align it with the ICRL setup by initializing task representations using randomly sampled trajectories and updating them after each episode.
@@ Appendix F Implementation Details of S-ICQL
Algorithm Hyperparameters. We evaluate the proposed S-ICQL algorithm on eight environments: Darkroom, Push, Reach, Cheetah‑Vel, Walker‑Param, Ant‑Dir, PickPlace and Humanoid-Dir. For all experiments, we use the Adam optimizer with a weight decay of 1e-4, gradient‑norm clipping at 10, an experience step of 4, and a batch size of 128. Table 10 summarizes the detailed hyperparameters of S-ICQL in each domain.
@@
Hyperparameters Darkroom Push Reach Cheetah-Vel Walker-Param Ant-Dir PickPlace Humanoid-Dir
Training steps $2\mathrm{e}5$ $4\mathrm{e}5$ $4\mathrm{e}5$ $4\mathrm{e}5$ $4\mathrm{e}5$ $4\mathrm{e}5$ $4\mathrm{e}5$ $8\mathrm{e}5$
Learning rate 3e-4 1e-4 1e-4 3e-4 3e-4 3e-4 1e-4 3e-4
Prompt horizon $h$ 100 100 100 200 200 200 200 200
Embedding dim 32 128 128 128 128 128 128 128
Output layers num 1 2 2 2 2 2 2 2
Temperature parameter $\lambda$ 0.01 0.001 0.001 0.001 0.1 0.01 0.01 0.1
Expectile parameter $\omega$ 0.7 0.5 0.5 0.7 0.7 0.7 0.5 0.7
Soft update $\alpha$ 0.005 0.005 0.005 0.005 0.005 0.005 0.005 0.005
Discount factor 0.99 0.99 0.99 0.99 0.99 0.99 0.99 0.99
Compute. We train our models on NVIDIA RTX3090 GPUs paired with an AMD EPYC 9654 CPU and 512GB of RAM. Pretraining the world model takes approximately 1–2 hours, while pretraining the causal transformer requires about 2–22 hours, depending on the environment’s complexity.
@@ Appendix G Hyperparameter Analysis
Appendix H Analysis of Loss Weight Coefficients
We conduct a hyperparameter analysis of Eq. 7 by evaluating multiple reweighting configurations for its three loss components, $\mathcal{L}_{\pi}(\theta)$, $\mathcal{L}_{Q}(\theta)$, and $\mathcal{L}_{V}(\theta)$. We vary the relative coefficients to emphasize different aspects of the learning objective, including configurations that up-weight the policy loss to prioritize policy fitting and configurations that increase the weight of the value or Q-value losses to strengthen critic learning, while keeping all other training settings fixed. As shown in Table 13, the performance degrades substantially when the policy loss is down-weighted (i.e., $\mathsf{c}_{1}=0.1$), while remaining insensitive to coefficients on value losses (i.e., $\mathsf{c}_{2}/\mathsf{c}_{3}=0.1$). The equal-weight configuration $(1,1,1)$ consistently offers a stable and straightforward choice, supporting the use of a balanced weighting scheme for the three loss components in our main experiments.
Table 13: Analysis of loss weight coefficients in S-ICQL.
Coefficients $(\mathsf{c}_{1},\mathsf{c}_{2},\mathsf{c}_{3})$ $(0.1,1,1)$ $(1,0.1,1)$ $(1,1,0.1)$ $(1,1,1)$
[…]
Reach 591.88 $\pm$ 88.81 789.29 $\pm$ 9.17 730.37 $\pm$38.46 806.97 $\pm$ 6.35